### v1

@@
[…]
Traveling purchaser problem; deep reinforcement learning; bipartite graph; policy network; meta-learning.
@@ 1 Introduction
The traveling purchaser problem (TPP) is a well-known combinatorial optimization (CO) problem, which has broad real-world applications and received much attention from both researchers and practitioners in recent decades (Burstall 1966, Goldbarg et al. 2009, Angelelli et al. 2016, Palomo-Martínez and Salazar-Aguilar 2019). In the TPP, given a list of products and their demand quantities, the purchaser aims to decide a route and an associated purchasing plan to meet the product demands by visiting a subset of markets, with the objective to minimize the sum of traveling and purchasing costs (Ramesh 1981).
Unfortunately, TPPs are known to be strongly NP-hard (Laporte et al. 2003). In particular, the main cha-
llenge in solving TPPs stems from the inherent coupling between routing and purchasing: the routing decisions have to take both traveling and purchasing costs into consideration, while the choice of visited markets will, in turn, impact the purchasing decisions and associated cost. Thus, existing works typically need to deal with both routing and purchasing simultaneously (Manerba et al. 2017), which, however, imposes many limitations. On the one hand, exact methods such as branch-and-cut (Laporte et al. 2003, Riera-Ledesma and González 2006) require joint optimization of routing and purchasing decisions, which generically leads to high computational cost, making them often intractable for realistically sized problem instances. For example, solving a medium-sized TPP instance with 100 markets and 50 products can take hours on a computer with a 2.3 GHz processor, and the computational time grows exponentially with the problem size (Riera-Ledesma 2012). On the other hand, various heuristic methods, such as the Generalized Savings Heuristic (GSH) (Golden et al. 1981), the Tour Reduction Heuristic (TRH) (Ong 1982), and the Commodity Adding Heuristic (CAH) (Pearn and Chien 1998), have been proposed to produce high-quality solutions for TPPs within a reasonable time. Readers are referred to Manerba et al. (2017) for further details. However, these heuristics have to carefully balance the effect of each operation on the traveling cost and purchasing cost, and thus rely heavily on sophisticated design that requires substantial human expertise and domain knowledge. Moreover, these hand-crafted heuristics typically need to be tailored case-by-case, resulting in that they may only be effective on limited instances with specific characteristics, and lack the ability to generalize across different instance distributions. For example, CAH performs poorly if the purchasing cost dominates in the optimization objective, and GSH may yield poor solutions if a market sells most of the products but is located far from other markets (Pearn and Chien 1998).
@@
In fact, the idea of leveraging DRL for CO problems is not entirely new though (Bengio et al. 2021, Cappart et al. 2022, Kim et al. 2024), particularly in routing problems such as the traveling salesman problem (TSP) (Vinyals et al. 2015, Bello et al. 2017, Kool et al. 2019, Kim et al. 2021) and the vehicle routing problem (VRP) (Nazari et al. 2018, Zhang et al. 2020, 2022, Xu et al. 2022). However, these efforts typically rely on problem-specific properties and are limited to problems with simple structures, which cannot be readily extended to TPPs. Therefore, considering the double nature of TPPs and building on our solution framework, we first propose a novel bipartite graph representation for TPPs, where the markets and products are represented as two sets of nodes, with the supply information (e.g., the supply quantities and prices) encoded as edge features between the market and product nodes. In contrast to existing representations for routing problems (e.g., the complete or k-NN graphs (Kool et al. 2019, Fu et al. 2021)), our bipartite graph representation can naturally capture the relations between markets and products through message-passing along edges. Then, we design a policy network, which can effectively extract information from the bipartite graph and use it for route construction. Specifically, our policy network exploits the connecting structure of the bipartite graph, and leverages graph neural networks (GNNs) and multi-head attention (MHA) to aggregate information both across and within the two sets of nodes. This effectively facilitates the extraction of relational information of the markets and products, such as the spatial relations between markets and the potential substitutions and complementarities in product supply, which is important for constructing a high-quality route. Moreover, it is worth mentioning that our bipartite graph representation and policy network have another appealing property: they are size-agnostic such that a trained policy network can be flexibly adapted to TPP instances of varying sizes without the need for retraining or parameter adjustment.
@@
In addition, despite our proposed framework and the key DRL components introduced above, how to efficiently learn a high-quality policy remains another challenge. Due to the huge state-action space, a randomly initialized policy network may suffer from inefficient exploration, which can result in slow convergence or even training collapse, particularly on large-sized instances. Different from the training techniques or tricks (Kwon et al. 2020, Kim et al. 2022) designed for specific problems (mainly TSP or VRP), we propose an effective and more generalizable training strategy based on meta-learning (Finn et al. 2017). The strategy trains an initialized policy network on a collection of small-sized instances, with the learning objective to achieve fast adaptation to large-sized instances at low fine-tuning cost. Furthermore, we observe that the initialized policy network can learn cross-distribution knowledge through meta-learning, such that it can effectively generalize across varied instance sizes and distributions, even achieving good zero-shot generalization performance to instances of much larger sizes that are never seen during training.
We empirically evaluate our DRL-based approach on two types of TPPs: the restricted TPP (R-TPP), where the supply quantities of available products at each market are limited, and the unrestricted TPP (U-
@@
TPP), which assumes unlimited supply quantities. We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark (Riera-Ledesma 2012). The results demonstrate that our DRL-based approach can produce near-optimal solutions with high computational efficiency, significantly outperforming well-established TPP heuristics in both solution quality and runtime. Notably, our approach achieves an average optimality gap consistently within 6% on each category of the TPPLIB benchmark in our experiment, yielding a reduction of 40%-90% compared to the baseline heuristics. In addition, we fur-
ther confirm the zero-shot generalization ability of the learned policy on instances that are much larger (up to 300 markets and 300 products, the largest size in TPPLIB) than the training instances.
@@ 2 Problem Formulation
In the TPP, a purchaser needs to buy a set of products from a set of markets. Each product can only be pur-
chased from certain markets, with potentially varying supply quantities and prices at different markets. The purchaser aims to decide a route that visits a subset of markets and an associated purchasing plan to meet all product demands, with the objective to minimize the sum of traveling and purchasing costs.
@@
[…]
@@
Then, the TPP can be formulated as follows (Laporte et al. 2003):
min $\quad\hskip-3.00003pt\sum_{(i,j)\in E}c_{ij}x_{ij}+\sum_{k\in K}\sum_{i\in M_{k}}p_{ik}z_{ik}$ (1)
@@
$\quad\hskip-12.74995pt\sum_{(i,j)\in\delta^{+}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (4)
$\quad\hskip-12.74995pt\sum_{(i,j)\in\delta^{-}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (5)
$\quad\hskip-12.74995pt\sum_{(i,j)\in\delta^{-}\left(M^{\prime}\right)}x_{ij}\geq y_{h},\;\forall M^{\prime}\subseteq M,h\in M^{\prime},$ (6)
@@
The MILP formulation (1)-(9) is commonly employed in exact methods for TPPs (Manerba et al. 2017). However, the number of constraints is exponential in the number of markets and the number of variables is also significantly large, which can lead to a massive branch-and-bound tree with weak lower bounds in an MILP solver. The branch-and-cut algorithms, as introduced in Laporte et al. (2003) and Riera-Ledesma and González (2006), attempt to alleviate this issue by dynamically generating variables and separating constraints, thereby to reduce the tree size and accelerate solving. Nonetheless, the dynamic pricing and separation procedures are still computationally expensive, making it often unaffordable to solve the MILP exactly in practice, especially for large-sized TPP instances. In contrast, our DRL-based approach does not directly solve the MILP, but instead creates a bipartite graph representation (Section 4) based on its structure, which is further used as input to the policy network for route construction.
@@
We now introduce our “solve separately, learn globally” framework for TPPs. As aforementioned, prior works typically require to address routing and purchasing simultaneously (Manerba et al. 2017). In contrast, by leveraging DRL, we can decouple the routing decisions and purchasing decisions at the operating level, i.e., “solve separately”, while guiding the decision-making using a policy, which is learned to optimize the global solution objective, i.e., “learn globally”. This framework offers the potential for both high computational efficiency and high-quality solutions if the policy is well-trained.
@@ 3.2 MDP Formulation
where $v^{1},v^{2},\dots,v^{t-1}$ denote the nodes selected at previous decision steps $1,2,\dots,t-1$, respectively, and $d_{k}^{t}$ denotes the remaining demand of product $k\in K$, assuming all possible quantities of product $k$ supplied in the partial route have been purchased. In the initial state $s_{1}$, the route starts from the depot $v_{0}$ and no products have been purchased, i.e., where $\pi^{1}=\left(v_{0}\right)$ and $d^{1}=\left(d_{1},d_{2},\dots,d_{\left|K\right|}\right)$.
@@
3.2.4 Reward.
After the sequential route construction process described above, the second stage, i.e.,
purchase planning, is executed, which is integrated into the reward calculation procedure for the terminal state. Given the constructed route, a linear transportation problem is solved to derive the optimal purchasing plan for the visited markets. In the transportation problem, each source point corresponds to a market on the route, and each destination point corresponds to a product to be purchased. The cost to purchase a unit of product $k$ from market $i$ is exactly the unit transportation cost from source point $i$ to destination point $k$ in the transportation problem. This transportation problem can be readily solved using an off-the-shelf LP solver or a polynomial-time dynamic programming algorithm (Munkres 1957). So far, a complete solution, including a route and the associated purchasing plan, has been obtained.
@@ 4 Bipartite Graph Representation for TPPs
Similar to existing DRL-based methods for routing problems (Kool et al. 2019, Fu et al. 2021), a natural and straightforward idea is to represent a TPP instance as a complete or k-NN graph, where each market is represented as a node, which is connected to each other, and the product supply information at each market are represented as node features. However, despite containing necessary information for the TPP instance, such representations impose many drawbacks when used in DRL. First, the feature dimensions vary with the number of products in the TPP instance being solved. For example, considering two features for each product: the available quantity and price, a 50-product instance requires a 100-dimensional feature vector for each node to record these 50 products, while an instance with 100 products would require 200 dimensions. This will lead to a computational mismatch in the forward-propagation of neural networks. In other words, a policy network can only be applied on fixed-sized TPP instances, unable to adapt to instances with different sizes even if only one additional product is included, which significantly limits its practical usability. Second, the relations between markets and products, such as the substitutions and complementarities in product supply between different markets, are very important for constructing a low-cost route, but such information is largely ignored in the aforementioned complete or k-NN graph representations. To address these limitations, we design a novel bipartite graph representation, which is built upon the MILP formulation of TPPs (Section 2).
Bipartite graphs have been demonstrated effective in representing general MILP problems (Nair et al. 2020, Morabit et al. 2021), where the variables and constraints are represented as two sets of nodes, and an
edge connects a variable node and a constraint node if the corresponding column contributes to this constraint. Readers are referred to Chen et al. (2023) for a more detailed introduction and analysis. However, directly applying this representation to the MILP formulation (1)-(9) for TPPs can introduce unnecessary redundancy: the parameters to characterize a specific TPP instance exist only in the objective function (1) and constraints (2)-(3), whereas constraints (4)-(6), which are used to form a simple cycle, have already been guaranteed by the masking rules (Section 3.2.2) and therefore need not be explicitly represented for the policy network. Moreover, the constraints and variables associated with the same element (e.g., a market or a product) can be further merged to reduce the graph size and thus improve the computational efficiency of the policy network.
@@
Therefore, we design a bipartite graph representation for TPPs based on the objective function (1) and constraints (2)-(3), with appropriate mergence for the constraints and variables therein. Our bipartite graph representation is shown as Figure 2. We represent the constraints associated only with products (i.e., constraints (2)) and the variables associated only with markets/depot (i.e., variables $v_{i}$ and $x_{ij}$) as two sets of nodes, and the constraints and variables associated with both products and markets (i.e., constraints (3) and variables $z_{ik}$) as edges between the product nodes and market/depot nodes. Specifically, each variable $v_{i}$ is represented as a market/depot node, and the objective coefficients $c_{ij}$ of variables $x_{ij}$ are represented as coordinate features of the market/depot nodes. Each constraint in (2) is represented as a product node, with the right-hand side values $d_{k}$ attached as node features. An edge connects a market node $i$ and a product node $k$ if the associated variable $z_{ik}$ is not always restricted to 0 in constraints (3), that is, product $k$ is available at market $i$. The edge feature is a 2-dimensional vector, which consists of the corresponding objective coefficient $p_{ik}$ of variable $z_{ik}$ and the constraint coefficient $q_{ik}$ in constraints (3). For simplicity in the remainder of this paper, we use the term “market nodes” to refer to “market/depot nodes” if it does not cause confusion. We remark that our bipartite graph representation can be viewed from a more intuitive perspective: the two core elements in TPPs, i.e., markets and products, are represented as two sets of nodes, and their connections, i.e., available quantities and prices of the products at the markets, are characterized using the edges between them.

### v5

@@ Abstract
keywords:
Traveling purchaser problem , deep reinforcement learning , bipartite graph , policy network , meta-learning
journal: arXiv
@@ 1 Introduction
The traveling purchaser problem (TPP) is a well-known combinatorial optimization (CO) problem, which has broad real-world applications and received much attention from both researchers and practitioners in recent decades (Burstall, 1966; Goldbarg et al., 2009; Angelelli et al., 2016; Palomo-Martínez and Salazar-Aguilar, 2019). In the TPP, given a list of products and their demand quantities, the purchaser aims to decide a route and an associated purchasing plan to meet the product demands by visiting a subset of markets, with the objective to minimize the sum of traveling and purchasing costs (Ramesh, 1981).
Unfortunately, TPPs are known to be strongly NP-hard (Laporte et al., 2003). In particular, the main challenge in solving TPPs stems from the inherent coupling between routing and purchasing: the routing decisions have to take both traveling and purchasing costs into consideration, while the choice of visited markets will, in turn, impact the purchasing decisions and associated cost. Thus, existing works typically need to deal with both routing and purchasing simultaneously (Manerba et al., 2017), which, however, imposes many limitations. On the one hand, exact methods such as branch-and-cut (Laporte et al., 2003; Riera-Ledesma and González, 2006) require joint optimization of routing and purchasing decisions, which generically leads to high computational cost, making them often intractable for realistically sized problem instances. For example, solving a medium-sized TPP instance with 100 markets and 50 products can take hours on a computer with a 2.3 GHz processor, and the computational time grows exponentially with the problem size (Riera-Ledesma, 2012). On the other hand, various heuristic methods, such as the Generalized Savings Heuristic (GSH) (Golden et al., 1981), the Tour Reduction Heuristic (TRH) (Ong, 1982), and the Commodity Adding Heuristic (CAH) (Pearn and Chien, 1998), have been proposed to produce high-quality solutions for TPPs within a reasonable time. Readers are referred to Manerba et al. (2017) for further details. However, these heuristics have to carefully balance the effect of each operation on the traveling cost and purchasing cost, and thus rely heavily on sophisticated design that requires substantial human expertise and domain knowledge. Moreover, these hand-crafted heuristics typically need to be tailored case-by-case, resulting in that they may only be effective on limited instances with specific characteristics, and lack the ability to generalize across different instance distributions. For example, CAH performs poorly if the purchasing cost dominates in the optimization objective, and GSH may yield poor solutions if a market sells most of the products but is located far from other markets (Pearn and Chien, 1998).
@@
In fact, the idea of leveraging DRL for CO problems is not entirely new though (Lopes Silva et al., 2019; Bengio et al., 2021; Que et al., 2023; Fajemisin et al., 2024), particularly in routing problems such as the traveling salesman problem (TSP) (Vinyals et al., 2015; Bello et al., 2017; Zhang et al., 2023; Gao et al., 2024) and the vehicle routing problem (VRP) (Nazari et al., 2018; Zhang et al., 2020; Kalatzantonakis et al., 2023; Rodríguez-Esparza et al., 2024). However, these efforts typically rely on problem-specific properties and are limited to problems with simple structures, which cannot be readily extended to TPPs. Therefore, considering the double nature of TPPs and building on our solution framework, we first propose a novel bipartite graph representation for TPPs, where the markets and products are represented as two sets of nodes, with the supply information (e.g., the supply quantities and prices) encoded as edge features between the market and product nodes. In contrast to existing representations for routing problems (e.g., the complete or k-NN graphs (Kool et al., 2019; Fu et al., 2021)), our bipartite graph representation can naturally capture the relations between markets and products through message-passing along edges. Then, we design a policy network, which can effectively extract information from the bipartite graph and use it for route construction. Specifically, our policy network exploits the connecting structure of the bipartite graph, and leverages graph neural networks (GNNs) and multi-head attention (MHA) to aggregate information both across and within the two sets of nodes. This effectively facilitates the extraction of relational information of the markets and products, such as the spatial relations between markets and the potential substitutions and complementarities in product supply, which is important for constructing a high-quality route. Moreover, it is worth mentioning that our bipartite graph representation and policy network have another appealing property: they are size-agnostic such that a trained policy network can be flexibly adapted to TPP instances of varying sizes without the need for retraining or parameter adjustment.
@@
In addition, despite our proposed framework and the key DRL components introduced above, how to efficiently learn a high-quality policy remains another challenge. Due to the huge state-action space, a randomly initialized policy network may suffer from inefficient exploration, which can result in slow convergence or even training collapse, particularly on large-sized instances. Different from the training techniques or tricks (Kwon et al., 2020; Kim et al., 2022) designed for specific problems (mainly TSP or VRP), we propose an effective and more generalizable training strategy based on meta-learning (Finn et al., 2017). The strategy trains an initialized policy network on a collection of small-sized instances, with the learning objective to achieve fast adaptation to large-sized instances at low fine-tuning cost. Furthermore, we observe that the initialized policy network can learn cross-distribution knowledge through meta-learning, such that it can effectively generalize across varied instance sizes and distributions, even achieving good zero-shot generalization performance to instances of much larger sizes that are never seen during training.
@@
We empirically evaluate our DRL-based approach on two types of TPPs: the restricted TPP (R-TPP), where the supply quantities of available products at each market are limited, and the unrestricted TPP (U-TPP), which assumes unlimited supply quantities. We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark (Riera-Ledesma, 2012). The results demonstrate that our DRL-based approach can produce near-optimal solutions with high computational efficiency, significantly outperforming well-established TPP heuristics in both solution quality and runtime. Notably, our approach achieves an average optimality gap consistently within 6% on each category of the TPPLIB benchmark in our experiment, yielding a reduction of 40%-90% compared to the baseline heuristics. In addition, we further confirm the zero-shot generalization ability of the learned policy on instances that are much larger (up to 300 markets and 300 products, the largest size in TPPLIB) than the training instances.
@@ 2 Problem Formulation
In the TPP, a purchaser needs to buy a set of products from a set of markets. Each product can only be purchased from certain markets, with potentially varying supply quantities and prices at different markets. The purchaser aims to decide a route that visits a subset of markets and an associated purchasing plan to meet all product demands, with the objective to minimize the sum of traveling and purchasing costs.
@@
The TPP can be formulated as a mixed integer linear programming (MILP) problem. Let $y_{i}$ be a binary variable taking value $1$ if market $i$ is selected, and 00 otherwise. Let $x_{ij}$ be a binary variable taking value $1$ if edge $(i,j)$ is traversed, and 00 otherwise. Let $z_{ik}$ be a variable representing the quantity of product $k$ purchased at market $i\in M_{k}$. For a subset of nodes $V^{\prime}\subset V$, we define:
@@
Then, the TPP can be formulated as follows (Laporte et al., 2003):
min $\quad\sum_{(i,j)\in E}c_{ij}x_{ij}+\sum_{k\in K}\sum_{i\in M_{k}}p_{ik}z_{ik}$ (1)
@@
$\quad\sum_{(i,j)\in\delta^{+}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (4)
$\quad\sum_{(i,j)\in\delta^{-}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (5)
$\quad\sum_{(i,j)\in\delta^{-}\left(M^{\prime}\right)}x_{ij}\geq y_{h},\;\forall M^{\prime}\subseteq M,h\in M^{\prime},$ (6)
@@
The MILP formulation (1)-(9) is commonly employed in exact methods for TPPs (Manerba et al., 2017). However, the number of constraints is exponential in the number of markets and the number of variables is also significantly large, which can lead to a massive branch-and-bound tree with weak lower bounds in an MILP solver. The branch-and-cut algorithms, as introduced in Laporte et al. (2003) and Riera-Ledesma and González (2006), attempt to alleviate this issue by dynamically generating variables and separating constraints, thereby to reduce the tree size and accelerate solving. Nonetheless, the dynamic pricing and separation procedures are still computationally expensive, making it often unaffordable to solve the MILP exactly in practice, especially for large-sized TPP instances. In contrast, our DRL-based approach does not directly solve the MILP, but instead creates a bipartite graph representation (Section 4) based on its structure, which is further used as input to the policy network for route construction.
@@
We now introduce our “solve separately, learn globally” framework for TPPs. As aforementioned, prior works typically require to address routing and purchasing simultaneously (Manerba et al., 2017). In contrast, by leveraging DRL, we can decouple the routing decisions and purchasing decisions at the operating level, i.e., “solve separately”, while guiding the decision-making using a policy, which is learned to optimize the global solution objective, i.e., “learn globally”. This framework offers the potential for both high computational efficiency and high-quality solutions if the policy is well-trained.
@@ 3.2.1 State
where $v^{1},v^{2},\dots,v^{t-1}$ denote the nodes selected at previous decision steps $11,2,\dots,t-$ 1, respectively, and $d_{k}^{t}$ denotes the remaining demand of product $k\in K$, assuming all possible quantities of product $k$ supplied in the partial route have been purchased. In the initial state $s_{1}$, the route starts from the depot $v_{0}$ and no products have been purchased, i.e., where $\pi^{1}=\left(v_{0}\right)$ and $d^{1}=\left(d_{1},d_{2},\dots,d_{\left|K\right|}\right)$.
@@ 3.2.3 Transition
3.2.4 Reward
After the sequential route construction process described above, the second stage, i.e., purchase planning, is executed, which is integrated into the reward calculation procedure for the terminal state. Given the constructed route, a linear transportation problem is solved to derive the optimal purchasing plan for the visited markets. In the transportation problem, each source point corresponds to a market on the route, and each destination point corresponds to a product to be purchased. The cost to purchase a unit of product $k$ from market $i$ is exactly the unit transportation cost from source point $i$ to destination point $k$ in the transportation problem. This transportation problem can be readily solved using an off-the-shelf LP solver or a polynomial-time dynamic programming algorithm (Munkres, 1957). So far, a complete solution, including a route and the associated purchasing plan, has been obtained.
@@ 4 Bipartite Graph Representation for TPPs
Similar to existing DRL-based methods for routing problems (Kool et al., 2019; Fu et al., 2021), a natural and straightforward idea is to represent a TPP instance as a complete or k-NN graph, where each market is represented as a node, which is connected to each other, and the product supply information at each market are represented as node features. However, despite containing necessary information for the TPP instance, such representations impose many drawbacks when used in DRL. First, the feature dimensions vary with the number of products in the TPP instance being solved. For example, considering two features for each product: the available quantity and price, a 50-product instance requires a 100-dimensional feature vector for each node to record these 50 products, while an instance with 100 products would require 200 dimensions. This will lead to a computational mismatch in the forward-propagation of neural networks. In other words, a policy network can only be applied on fixed-sized TPP instances, unable to adapt to instances with different sizes even if only one additional product is included, which significantly limits its practical usability. Second, the relations between markets and products, such as the substitutions and complementarities in product supply between different markets, are very important for constructing a low-cost route, but such information is largely ignored in the aforementioned complete or k-NN graph representations. To address these limitations, we design a novel bipartite graph representation, which is built upon the MILP formulation of TPPs (Section 2).
Bipartite graphs have been demonstrated effective in representing general MILP problems (Nair et al., 2020; Morabit et al., 2021), where the variables and constraints are represented as two sets of nodes, and an edge connects a variable node and a constraint node if the corresponding column contributes to this constraint. Readers are referred to Chen et al. (2023) for a more detailed introduction and analysis. However, directly applying this representation to the MILP formulation (1)-(9) for TPPs can introduce unnecessary redundancy: the parameters to characterize a specific TPP instance exist only in the objective function (1) and constraints (2)-(3), whereas constraints (4)-(6), which are used to form a simple cycle, have already been guaranteed by the masking rules (Section 3.2.2) and therefore need not be explicitly represented for the policy network. Moreover, the constraints and variables associated with the same element (e.g., a market or a product) can be further merged to reduce the graph size and thus improve the computational efficiency of the policy network.
@@
Therefore, we design a bipartite graph representation for TPPs based on the objective function (1) and constraints (2)-(3), with appropriate mergence for the constraints and variables therein. Our bipartite graph representation is shown as Figure 2. We represent the constraints associated only with products (i.e., constraints (2)) and the variables associated only with markets/depot (i.e., variables $v_{i}$ and $x_{ij}$) as two sets of nodes, and the constraints and variables associated with both products and markets (i.e., constraints (3) and variables $z_{ik}$) as edges between the product nodes and market/depot nodes. Specifically, each variable $v_{i}$ is represented as a market/depot node, and the objective coefficients $c_{ij}$ of variables $x_{ij}$ are represented as coordinate features of the market/depot nodes. Each constraint in (2) is represented as a product node, with the right-hand side values $d_{k}$ attached as node features. An edge connects a market node $i$ and a product node $k$ if the associated variable $z_{ik}$ is not always restricted to 00 in constraints (3), that is, product $k$ is available at market $i$. The edge feature is a 2-dimensional vector, which consists of the corresponding objective coefficient $p_{ik}$ of variable $z_{ik}$ and the constraint coefficient $q_{ik}$ in constraints (3). For simplicity in the remainder of this paper, we use the term “market nodes” to refer to “market/depot nodes” if it does not cause confusion. We remark that our bipartite graph representation can be viewed from a more intuitive perspective: the two core elements in TPPs, i.e., markets and products, are represented as two sets of nodes, and their connections, i.e., available quantities and prices of the products at the markets, are characterized using the edges between them.